OpenStudy (kittiwitti1):

If \[\int_{1}^{8}f(x)dx=9.6\] and \[\int_{5}^{8}f(x)dx=2.5\] then find \[\int_{1}^{5}f(x)dx\]

OpenStudy (irishboy123):

just add them up ..... make sense?? this is about the linearity of integration.

OpenStudy (kittiwitti1):

i hate that the latex resets when people comment... what do you mean add up? o_o

OpenStudy (irishboy123):

are you online right now?

OpenStudy (irishboy123):

cos @sshayer is just gonna type out the answer !!!!!

OpenStudy (sshayer):

\[\int\limits_{a}^{b}f(x)dx=\int\limits_{a}^{c}f(x)dx+\int\limits_{c}^{b}f(x)dx\]

OpenStudy (sshayer):

whenever he is on line ,he will read the formula

OpenStudy (irishboy123):

@sshayer this site is apparently dying but your stuff is brilliant! good health, mate !!

OpenStudy (sshayer):

thank you.

OpenStudy (irishboy123):

mp, mate :-)

OpenStudy (kittiwitti1):

I lagged out... sorry... and uh... I'm female @sshayer o_O

OpenStudy (kittiwitti1):

ok... should I add this by \(\Sigma\) summation or something?

OpenStudy (irishboy123):

no! this: \[\int\limits_{a}^{b}f(x)dx=\int\limits_{a}^{c}f(x)dx+\int\limits_{c}^{b}f(x)dx\]

OpenStudy (kittiwitti1):

I don't get how to add that

OpenStudy (kittiwitti1):

should I convert to summation or ?

OpenStudy (irishboy123):

@kittiwitti1 \(9.6 - 2.5 = ?\)

OpenStudy (kittiwitti1):

how did you come about those numbers? o_o

OpenStudy (kittiwitti1):

7.1

OpenStudy (irishboy123):

your orig post: \[\int_{1}^{8}f(x)dx=9.6\] and \[\int_{5}^{8}f(x)dx=2.5\] then find \[\int_{1}^{5}f(x)dx\]

OpenStudy (kittiwitti1):

ahh... I see sorry, LaTeX glitched back to code once a comment appeared

OpenStudy (irishboy123):

but @sshayer is spesh !!

OpenStudy (kittiwitti1):

is there a way to add it when no values for the integral are given? maybe but he thought I was male... lol :p

OpenStudy (irishboy123):

fo !! you're now having a laugh

OpenStudy (irishboy123):

:-))

OpenStudy (kittiwitti1):

er... got confused. what do you mean? :(

OpenStudy (irishboy123):

nothing are we moving ahead?

OpenStudy (sshayer):

\[\int\limits_{1}^{8}f(x)dx=\int\limits_{1}^{5}f(x)dx+\int\limits_{5}^{8}f(x)dx\]

OpenStudy (sshayer):

sorry,kittywitti1 i thought you are male.

OpenStudy (kittiwitti1):

ah... it is ok I guess. was just a mistake @sshayer I forgive lol

OpenStudy (irishboy123):

wtf?!?!

OpenStudy (kittiwitti1):

what?

OpenStudy (kittiwitti1):

something I missed?

OpenStudy (irishboy123):

nah, my bad

OpenStudy (kittiwitti1):

ok. lol I will close this question now... can I have a medal though? :(

OpenStudy (irishboy123):

i gave your medal to ssh he declined but i insisted!! seriously :)

OpenStudy (kittiwitti1):

uh... ok? :(